OpenStudy (anonymous):

Help please 4^3x4^12 4 is bold (4^12) 46^36 4^15 4^12 4^9

OpenStudy (butterfield1215):

whats 12+3

OpenStudy (anonymous):

15

OpenStudy (butterfield1215):

when multiplying you will add the roots together when dividing you will subtract them

OpenStudy (anonymous):

so 4 ^15

OpenStudy (butterfield1215):

yep

OpenStudy (anonymous):

ty

OpenStudy (butterfield1215):

No problem

OpenStudy (butterfield1215):

do you need help with anything else if so just pm me

OpenStudy (anonymous):

pm u

OpenStudy (butterfield1215):

whats the question

OpenStudy (anonymous):

3^3 x 3^33 3^24 3^30 3^36 3^99

OpenStudy (butterfield1215):

ok agian your multiply so you?

OpenStudy (anonymous):

3+33=36

OpenStudy (butterfield1215):

good job

OpenStudy (anonymous):

oh so 3^36 lol

OpenStudy (butterfield1215):

Also when doing this the bottom numbers have to be the same

OpenStudy (anonymous):

i look dumb right now

OpenStudy (butterfield1215):

No you dont

OpenStudy (anonymous):

oh ok :)

OpenStudy (anonymous):

can you help me with 3 more ?

OpenStudy (butterfield1215):

yep

OpenStudy (anonymous):

i promise ill give you medals

OpenStudy (anonymous):

Suppose you earned 8t - 3 dollars on Monday and 6t + 5 dollars on Tuesday. What were you total earnings? Simplify your answer. A) 2t + 2 B) 2t - 8 C) 14t + 2 D) 14t - 8

OpenStudy (butterfield1215):

@ortiz09

OpenStudy (butterfield1215):

I need some clarification on this one

OpenStudy (anonymous):

sorry worng ? Suppose you earned 6t +2 dollars on Monday and 9t -6 dollars on Tuesday. What were you total earnings? Simplify your answer. A)15t+ 4 B) 15t +8 C) 3t - 4 D) -3t+8

OpenStudy (butterfield1215):

oh ok

OpenStudy (anonymous):

it looks like d

OpenStudy (butterfield1215):

Yes it does because i could only see you setting them to eachother but im not 100% on this one

OpenStudy (anonymous):

ok ty and 2 more

OpenStudy (butterfield1215):

ok post them both at the same time

OpenStudy (anonymous):

22. A.write an expression in simplest form for the perimeter of each triangle B.write another expression in simplest form that shows the difference between the perimeter of the large triangle and the perimeter of the small triangle

OpenStudy (butterfield1215):

what do i need to

OpenStudy (butterfield1215):

oh nevermind

OpenStudy (butterfield1215):

@sleepyjess

OpenStudy (butterfield1215):

I dont really know this one because its been so long since i have done it

OpenStudy (anonymous):

oh ok i figure it out then ill try

OpenStudy (butterfield1215):

brb

OpenStudy (anonymous):

but can you give me a hint on how to find the answer

OpenStudy (anonymous):

ok

OpenStudy (butterfield1215):

wheres the table at

OpenStudy (anonymous):

I found out who was right it was e something for 23

OpenStudy (butterfield1215):

ok

OpenStudy (anonymous):

can you check my answer for 24

OpenStudy (butterfield1215):

yeah

OpenStudy (butterfield1215):

brb

OpenStudy (butterfield1215):

ok where is number 24

OpenStudy (anonymous):

yes because 11^0=1 and11^211^2 =11^2 11^2=11*1111*11=1

OpenStudy (butterfield1215):

wait what

OpenStudy (anonymous):

so here what it says evaluate the following expression 11^0 and 11^2/11^2 are the answer to parts the same explain why or why not

OpenStudy (butterfield1215):

anthing to the 0 power is 1

OpenStudy (anonymous):

oki thanks im going to tag you 3 times ok so you can get a medal :)

OpenStudy (butterfield1215):

ok